- A15/22
- B7/22
- C5/33
- D5/12
Step-by-Step Solution
Total balls = 12, so total ways = C(12,2) = 66. It is easier to use the complement: 'no red' means both come from the 7 non-red balls, C(7,2) = 21, giving probability 21/66 = 7/22. Hence P(at least one red) = 1 - 7/22 = 15/22.
Answer: A) 15/22